# FutureGen: LLM-RAG Approach to Generate the Future Work of Scientific Article

### v1

@@ 6.3 Evaluation of Generation
Evaluation-incorporating LLM feedback. After generating the ‘future work’ text, a separate LLM served as a judge, scoring the output on a 1–5 scale with justification (feedback). This feedback was then integrated into the prompt to refine the generated text further. We evaluated the LLM-generated text after each iteration with GPT GT, and our results show a significant performance improvement across all metrics after incorporating LLM feedback once (Table 1). Moreover, from Table 3, we observed that after the first feedback loop (iteration 2), using GPT-4o mini + RAG, performance improved across nearly all metrics. However, in iteration 3, performance declined when the model received feedback a second time, indicating that excessive iterations introduced bias. These findings suggest that a single round of feedback integration is optimal for improving model performance without introducing unwanted biases.
@@
Evaluation-among all models. We applied various models, and among all of the models, GPT 4o mini with RAG performs best when LLM feedback goes to the model (iteration 2) when considering ‘GPT Refined Ground Truth’ in NLP-based metrics (Table 3) and metrics in LLM as a Judge (Table 2). In terms of readability, Llama 3 zero shot performs better. We also compared our model (GPT 4o mini + RAG) with the simple zero shot model of GPT 4o mini, and we observed gains in Jaccard Similarity (+1.08), BLEU score (+0.7), and, most notably, BERTScore F1 (+0.23) (Table 3). However, n-gram-based metrics like ROUGE-L (-3.5) decreased due to the introduction of new information from RAG, which diluted exact text overlap.
6.4 Human Evaluation
Evaluating with only LLM-based evaluation metrics raises the question of reliability and potential biases. To overcome this problem, we compared LLM-based evaluation with human judgments and found a strong relationship. We conducted a user study with 200 randomly selected samples involving three annotators who rated below two questions from 1 to 3 (1 = worst, 3 = best). The annotators are graduate students, knowledgeable about machine learning, not related with this paper. The questions are: Q1: How well does LLM extract the future work (ground truth) without generating anything? Q2: How good has LLM generated future work (from Abstract, Introduction, and Conclusion) based on ground truth (paper mentioned future work)? Q3: How does the LLM Feedback approach improve performance in terms of originality? In Table 5, the average rating across all questions is approximately 2, reflecting an overall average performance. The rating for Q1 suggests that the LLM is effective at extracting text. Comparing LLM-generated future work with ground truth (Q2) shows average performance supported by strongly kappa and Kendall’s tau scores. Notably, in Q3, our proposed approach with one feedback iteration achieves an average human rating of 2.75, demonstrating that the LLM feedback loop significantly enhances result quality (Details A.1).
6.5 Trends in NLP Future Work
We analyzed the evolution of future work directions in NLP from 2012 to 2024 (Table 9, 7, Appendix). The early years emphasized neural machine translation, model improvement, and basic future work identification. From 2017 to 2019, there was a strong focus on language modeling, machine translation, and multilingual text representation, reflecting the rise of deep learning in NLP. From 2020 to 2021, research has expanded towards multi-scale learning techniques and generalization. In recent years, evaluation frameworks, bias mitigation, and performance robustness have increasingly been prioritized, highlighting the community’s shift toward more reliable, explainable, and fair NLP models. Early research focused on statistical methods, machine translation, and model improvement. Still, recent trends emphasize LLMs, prompt engineering, and data efficiency, also increasingly emphasize model performance robustness, privacy, adversarial attacks, generalization, and evaluation in NLP models (Figure 2, Table 14, 7 Appendix). From the topic analysis from the recent year (ACL 2024), we can see that the prominent topics are ‘linguistic,’ ‘LLM,’ ‘privacy,’ ‘annotation,’ etc. (Table 12, Appendix). Researchers also emphasize the need for better evaluation methods to assess and mitigate model biases, fairness, and societal implications of NLP models. Ensuring NLP models are trustworthy, explainable, and interpretable evaluation frameworks for document generation and annotation quality. Over time, the focus has shifted from a narrow concentration on model and translation improvements toward more nuanced themes that encompass ethical considerations, interdisciplinary approaches, and real-world applicability (Table 13, Appendix). In addition, Future research increasingly explore connections of interdisciplinary approaches between AI and human cognition.
@@ 6.5 Trends in NLP Future Work
Q2 1.92 0.30 0.36
@@ 7 Discussions
Our experiment shows that GPT-Refined Ground Truth aligns well with LLM-generated future work text, indicating that RAW Ground Truth contains more noise and irrelevant sentences. Our proposed approach can extract future work-related sentences from papers that lack explicit future work sections. Also, incorporating RAG with GPT 4o mini increased the Jaccard Similarity and BERTScore but reduced the n-gram overlaps matching than zero shot GPT 4o mini. Notably, one-time LLM feedback improves the performance of LLM when considering GPT Refined Ground Truth. To make generalizability, our model showed incorporating a single round of feedback led to significant improvements. However, applying a second round of feedback negatively impacted performance, producing a more biased response. We compared using all sections versus only the top three sections as input. Although restricting our analysis to the top three sections resulted in a minor performance drop (Table 4), we adopted this approach to mitigate the computational and API costs. The feedback might miss issues related to the originality of ideas or the soundness of the proposed future work. To alleviate this problem, we conducted a human experiment (Q3) and got a good average rating (2.75). We relied on the same LLM (GPT 4o mini) as the generator and evaluator, which enhanced the LLM’s own bias, we experimented with Llama 3 70b as an evaluator, but Llama provided unncessary higher scores even though sentence quality was not good (Table 11, Appendix).
@@ 8 Conclusions
The future work section is a forward-looking guide, helping the research community explore new directions. We selected the top three sections from each research paper to mitigate LLM context size limitations. We then utilized an LLM to extract future work, producing a more coherent ground truth that enhances model performance. Additionally, we integrated an external vector database to further improve LLM’s performance to generate future work from input text. For evaluation, we applied NLP-based metrics alongside an LLM-as-a-Judge approach, using explainable LLM metrics to assess performance, provide feedback, and iteratively refine text generation. Furthermore, we conducted a 13-year trend analysis to examine how future work priorities have evolved over time, highlighting key shifts in research focus.
@@
The study focuses on ACL papers (2012–2024), providing strong domain relevance but limiting cross-disciplinary applicability. RAG retrieval was limited to 100 papers due to computational costs. Employing the same LLM as both generator and evaluator may introduce systematic biases that affect the output scores. Moreover, we assumed that the LLM-extracted Future Work could serve as the ground truth, that incorporating LLM feedback enhances performance, and that the suggestive Future Work generated by the LLM aligns with the ground truth—even though these evaluations were based on a limited number of annotators and samples. While the LLM evaluates coherence, reliability, readability, and grammatical correctness, these criteria may not fully capture essential academic qualities such as the originality of ideas or the robustness of the proposed future work and leads to overfitting. Furthermore, iterative refinement using the same model risks converging towards a narrow stylistic range and set of ideas, potentially increasing the likelihood of hallucinated content. For T5 and BART, we limited input length to 512 and 1024 tokens, respectively, which may restrict the model’s ability to capture complex research contexts fully. Also, we did not explore LLM feedback with Chain of Thought or self-consistency approaches, nor did we analyze hallucination levels with and without RAG integration. In the future, we will expand to broader research fields and refine extraction methods for implicit future work mentions. We will experiment on open-source LLMs to reduce API costs and conduct large-scale experiments to construct a comprehensive vector database to enhance retrieval quality in RAG. We will experiment with more human annotators to check whether the LLM as an extractor, generator, and iterative feedback loop sufficiently addresses deeper qualitative aspects of academic writing in the context of real-world applicability and make gold standard dataset. Moreover, to alleviate the bias from LLM feedback, we will experiment with Reinforcement Learning from Human Feedback (RLHF) and the AI Feedback (RLAIF) approach and make a more robust model. Also, we will investigate advanced LLM reasoning techniques, such as Chain of Thought and self-consistency, to refine LLM feedback mechanisms and expand future work extraction by incorporating cited papers and reviewers’ perspectives to diversify insights.
@@ A.1 Human Evaluation
Figure 3: BERTopic + LLM for generating title in future work.

### v2

@@ 6.3 Future Work Generation
Evaluation-incorporating LLM feedback. After generating the ‘ Future Work’ text, a separate LLM served as a judge, scoring the output on a 1–5 scale with justification (feedback). This feedback was then integrated into the prompt to refine the generated text further. We evaluated the LLM-generated text after each iteration with GPT GT, and our results show a significant performance improvement across all metrics after incorporating LLM feedback once (Table 3). Moreover, from Table 6, we observed that after the first feedback loop (iteration 2), using GPT-4o mini + RAG, performance improved across nearly all metrics. However, in iteration 3, performance declined when the model received feedback a second time, indicating that excessive iterations introduced bias. These findings suggest that a single round of feedback integration is optimal for improving model performance without introducing unwanted biases. In Table 4, using the GPT 4o mini model, we added self-feedback in a second iteration in NeurIPS data, excluding the NLP metrics, other LLM-based metrics such as coherence, relevance, and novelty improved the results. In the first iteration, the novelty score was 7.28, indicating a strong similarity between the LLM-generated future work and the ground truth. In the second iteration, we incorporated self-feedback, which helped uncover more underexplored ideas across multiple directions and resulted in increased +0.72 (Table 4). In GPT 4o mini + RAG settings, all LLM-based metrics exceeded the threshold of 3 on the combined Future Work + OpenReview (FW + OR) ground-truth set, so we did not apply additional feedback in this case.
@@
Evaluation-among all models. We applied various models, and among all of the models, GPT 4o mini with RAG performs best when LLM feedback goes to the model (iteration 2) when considering ‘GPT extracted Ground Truth’ in NLP-based metrics (Table 6) and LLM based metrics (Table 5). GPT 4o mini + RAG achieved better results than zero-shot GPT 4o mini, except n-gram based methods (Rouge-1,2, L) due to the introduction of new information from RAG (Table 6). In NeurIPS data (Table 4), incorporating RAG with GPT 4o mini didn’t increase results in NLP metrics, but it boosts performance on some LLM-based metrics.
Evaluation incorporating OpenReview We evaluated our approach on NeurIPS data using three ground-truth sources: Author-Mentioned (FW), OpenReview (OR), and their combination (FW+OR), as shown in Table 4. Incorporating OpenReview feedback (FW + OR) in ground truth yielded improvements across all metrics compared to other ground truths (FW, OR) in both models GPT 4o mini and GPT 4o mini + RAG. Adding OpenReview (OR) with future work in ground truth achieves lower novelty, meaning most LLM-generated ideas are covered.
LLM as a Judge We assessed the generated suggestions against the ground truth using six criteria—coherence, relevance, readability, grammar, overall impression, and novelty—by treating an LLM as the evaluator. Our primary judge was GPT-4o-mini, and to mitigate any generator–evaluator bias, we also tried Llama 3 70B. However, we found that Llama 3 tended to assign overly high scores even to low-quality outputs.
Hallucination Rate We evaluated hallucinations by treating each LLM-generated text as the hypothesis and the concatenation of the original input text and ground-truth future work as the premise. For each pair, we asked GPT-4o-mini to classify the relationship as entailment, neutral, or contradiction. Using RAG, GPT-4o-mini’s hallucination rate dropped to 19.52%, compared to the higher rate observed without RAG (Table 8).
Model Hallucincation Rate
GPT 4o mini 26.26 %
GPT 4o mini+RAG 6.74 %
Table 8: Hallucination rate of each model in NeurIPS data (lower is better)
@@
Q2 2.12 0.30 0.36
Q3 2.75 0.28 0.29
Table 9: Average rating and annotators user agreement on user study.
6.4 Human Evaluation
Using the same LLM as extractor, generator, and evaluator may raise potential biases such as self-validation bias and confirmation bias. To overcome this problem, we compared LLM-based evaluation with human judgments and found a strong relationship. We conducted these questions with a human: Q2: How good is LLM’s generated Future Work based on the ground truth? Q3: How does the LLM Feedback approach improve performance in terms of originality? In Table 11, the average rating across all questions is more than 2, reflecting an overall average performance. Comparing LLM-generated Future Work with ground truth (Q2) shows average performance (2.12) supported by strong kappa and Kendall’s tau scores. Notably, in Q3, our proposed approach with one feedback iteration achieves an average human rating of 2.75, demonstrating that the LLM feedback loop significantly enhances result quality.
Future Work
Future Research in Natural Language Processing and Model..
Future Work in Natural Language Processing and Image Captioning
Future Directions in NLP Model Development and Sentiment Analysis
Sentiment Analysis and Dispute Detection Models in Conversational
Future Directions in Language Modeling and Machine Translation
Future Research Directions in NLP Models and Techniques
Future Directions in Model Development for Multilingual..
Neural Machine Translation and Future Research Directions in NLP
Future Research Directions in NLP Models and Methods
Future Research Directions in Neural Machine Translation and ..
Evaluation Frameworks and Future Work in Generalization and ..
Future Directions in NLP Model Evaluation and Bias
Future Directions in Language Model Research and Performance ..
Table 10: Prominent Future Work in each year (2012-2024) in ACL data.
6.5 Trends in NLP Future Work
We analyzed the evolution of future work directions in NLP from 2012 to 2024 (Table 15, 13, Appendix). The early years emphasized neural machine translation, model improvement, and basic future work identification. From 2017 to 2019, there was a strong focus on language modeling, machine translation, and multilingual text representation, reflecting the rise of deep learning in NLP. From 2020 to 2021, research has expanded towards multi-scale learning techniques and generalization. In recent years, evaluation frameworks, bias mitigation, and performance robustness have increasingly been prioritized, highlighting the community’s shift toward more reliable, explainable, and fair NLP models. Early research focused on statistical methods, machine translation, and model improvement. Still, recent trends emphasize LLMs, prompt engineering, and data efficiency, also increasingly emphasize model performance robustness, privacy, adversarial attacks, generalization, and evaluation in NLP models (Figure 3, Table 20, 13 Appendix). From the topic analysis from the recent year (ACL 2024), we can see that the prominent topics are ‘linguistic,’ ‘LLM,’ ‘privacy,’ ‘annotation,’ etc. (Table 18, Appendix). Researchers also emphasize the need for better evaluation methods to assess and mitigate model biases, fairness, and societal implications of NLP models. Ensuring NLP models are trustworthy, explainable, and interpretable evaluation frameworks for document generation and annotation quality. Over time, the focus has shifted from a narrow concentration on model and translation improvements toward more nuanced themes that encompass ethical considerations, interdisciplinary approaches, and real-world applicability (Table 19, Appendix). In addition, Future research increasingly explore connections of interdisciplinary approaches between AI and human cognition.
Ques. Rating (Avg.) W. Kappa Kend’ Tau
Q1 2.54 0.24 0.29
Q2 2.12 0.30 0.36
@@ 7 Discussions
Our framework is readily extensible to other fields, such as biomedicine or social sciences, by harnessing peer-review feedback to build a robust ground truth, constructing domain-specific retrieval corpora, applying LLM-based evaluation metrics (including novelty assessments), and iteratively refining outputs via LLM self-feedback. Moreover, our experiment shows that GPT-extracted Ground Truth aligns well with LLM-generated Future Work text, indicating that the Tool extracted Ground Truth contains more noise and irrelevant sentences. Our proposed approach can extract Future Work-related sentences from papers that lack explicit Future Work sections. Also, incorporating RAG with GPT-4o mini increased performance in most of the metrics, except the n-gram overlaps and reducing hallucinations. Notably, one-time LLM feedback improves the performance of LLM when considering the GPT extracted Ground Truth. To make generalizability, our model showed that incorporating a single round of feedback led to significant improvements. However, applying a second round of feedback negatively impacted performance, producing a more biased response. Also, considering the top 3 sections instead of all sections results in minor performance drops. Moreover, we incorporated long-term goal suggestions from peer review from OpenReview to make a robust ground truth, which increased the performance in every metric. LLM shows strong relations with human evaluators in the extractor, generator, and feedback provider tasks.
@@ 8 Conclusions
The Future Work section is a forward-looking guide, helping the research community explore new directions. We utilized an LLM to extract Future Work, producing a more coherent ground truth that enhances model performancea and incorporated a strong ground truth from OpenReview. Additionally, we integrated an external vector database to further improve LLM’s performance to generate Future Work from input text. For evaluation, we applied NLP-based metrics alongside an LLM-as-a-Judge approach, using explainable LLM metrics to assess performance, provide feedback, and iteratively refine text generation. Furthermore, we conducted a 13-year trend analysis to examine how Future Work priorities have evolved over time, highlighting key shifts in research focus.
@@
Our analysis is confined to ACL papers (2012–2024) and NeurIPS (2021-22), which ensures domain relevance but limits cross-disciplinary generality. RAG retrieval was capped at 100 papers for cost reasons, and using the same LLM as both generator and evaluator may introduce bias. We treated LLM-extracted future work as ground truth and assumed that LLM-driven feedback and suggestions align with that standard, despite relying on a small pool of annotators and samples. Our evaluation criteria, coherence, relevance, readability, and grammar, do not fully capture scholarly attributes like originality or robustness, and iterative refinement with a single model risks stylistic convergence and potential hallucination. We limited T5 and BART inputs to 512 and 1,024 tokens and did not explore advanced feedback strategies (e.g., chain-of-thought, self-consistency). Taking 100 random papers from the RAG database can raise issues, such as generating future work of a paper where the data comes from after the publication date of the paper. To alleviate this problem, in the future we will take the related papers which is published before the actual paper. Also, in the future, we will extend our pipeline to additional research domains and improve methods for extracting implicit future-work mentions, evaluate open-source LLMs to reduce API costs, and build a large-scale domain-specific vector store to enhance RAG retrieval. We also plan to involve more human annotators to validate our extractor–generator–feedback loop and create a gold-standard dataset, mitigate evaluator bias through RLHF and RLAIF, and integrate advanced reasoning techniques while incorporating cited literature and reviewer perspectives to diversify and enrich future-work suggestions.